OpenStudy (anonymous):

Probability: If A and B are mutually exclusive events, P(A)=.26, and P(B)=.45, find (a) P(A union B), (b) , (d) P(not A union not B).

OpenStudy (anonymous):

(Since there is no way to put a "-" above a letter I'm going to define "prime" as not instead). Let me rewrite the problems: (a) \[P(A U B)\] (b)\[P(A \prime U B \prime)\] The probability rule of complement is:\[P(A\prime)=1-P(A)\]

OpenStudy (anonymous):

Do u know what are mutually exclusive events?

OpenStudy (anonymous):

Since they are mutually exclusive events: P(A or B) =P(A) + P(B)

OpenStudy (anonymous):

I do.

OpenStudy (anonymous):

Both events when added have to be either less than or equal to 1, not greater than 1, and contain no negative numbers.

OpenStudy (anonymous):

From my book, the first question is actually \[P(A \prime)\] and by using the probability rule of complement: P(not A)=1-P(A), I get: 1-.26 = .45 .45+.26=.71